Referring to problem 2 above. City A has 8,000 households and on average each household spends $50 per year on lighting fixtures and accessories. City B has 13,000 households, who on average are willing to spend $55 per year on lighting fixtures and accessories. City C has 18,000 households, who are willing to spend an average of $60 per year on lighting fixtures and accessories. As an electrical supplier, you have stores in each city and have allocated space 100 square feet dedicated to displaying lighting fixtures and accessories in each store. The electrical supplier plans to create a master lighting fixtures and accessories showroom in one of his existing stores, by expanding the space allocated to lighting fixtures and accessories in one store from 100 square feet to 400 square feet. Based upon the retail saturation index, which city would you suggest to the electrical supplier to make the expansion?
Answer #1

Retail saturation index is defined as below

IRS = Demand / Current retail selling space= (Population in area ;likely to buy products * per capita retail expenditure in area on products)/ Current retail space in sq. ft. in area selling products

IRS (City A) = (8000 * $50)/ 100 = 4,000

IRS (City B) = (13000 * $55)/ 100 = 7,150

IRS (City C) = (18000 * $60)/ 100 = 10,800

Based on retail saturation index (RSI), electrical supplier should make an expansion in City C (with high RSI).